OpenStudy (anonymous):

Solve for x on 0 =< x < 2pi : sin^3(2x) - sin2x= 0.

OpenStudy (***[isuru]***):

hi, \[\sin ^{3}(2x) - \sin (2x) = 0\] \[\sin ^{3}(2x) = \sin (2x)\] \[\sin ^{2} (2x) = 1\]\[\sin (2x) = \pm \sqrt{1} =\pm1\] when sin(2x) = +1\[\sin (2x) = \sin \frac{ \pi }{ 2 }\] the formula for a sin angle is...\[\theta = n \pi + (-1)^{n} \alpha \ ; \ ( n \ \in \ \mathbb{Z} ) \] which means.. \[2x = n \pi + (-1) ^{n}\frac{ \pi }{2 }\ ( n \ \in \ \mathbb{Z} )\]\[x =\frac{ n \pi }{ 2 } + (-1)^{n}\frac{ \pi }{ 4 }\] so the values for x are... \[x = \left\{ \frac{ \pi }{ 4 } , \frac{ 9\pi }{ 4 } , \frac{ 5 \pi }{4 } , \right\}\] did u get it so far?

OpenStudy (***[isuru]***):

now for the sin (2x) = -1 \[\sin (2x) = \sin (-\frac{ \pi }{ 2 })\] which means ... according to the above formula.. \[2x = n \pi + (-1)^{n} (-\frac{ \pi }{ 2 })\]\[x = \frac{ n \pi }{2 } - (-1)^{n}\frac{ \pi }{4 }\] so... the values for "x" r.. \[x = \left\{ \frac{ 3 \pi }{4 } , \frac{ 7 \pi }{4 } \right\}\] hope this will help ya!!

OpenStudy (anonymous):

@***[ISURU]*** Thank you!! :)